Suppose Jim has a demand curve of: and Sam has a demand curve of: oS 6 0.5p. How much is the total demand at a price of $9.00? 1.5 unit(s) (round your answer to two decimal places) How much is total demand at a price of $7.50? units (round your answer to two decimal places).
Answer #1

Writing the inverse demand functions

For Jim : p = 8-Q

For Sam : .5p = 6-Q

P=12-2*Q

Thus the maximim willingness to pay for Jim is $8, so at price of 9, demand comes from only Sam , hence total demand is Q = 6-.5*9

= 1.5 units

now at price = 7.5, both have a positive demand

So total demand is sum of two individual demands = .5 + 2.25

( 8-.5 + 6-.5*7.5) = .5+2.25 = 2.75 units